# revtex4-2Repair the float

### v1

@@ Abstract
The plots of Figure 6 show that incoherent Pauli errors have, as expected, a significant influence on the QCNN output for the Toric Code ground states as soon as the error rates reach the threshold $p_{\mathrm{th}}$. Up to the threshold, the QCNN output behaves similarly as in Figure 3. However, for $p_{X}=p_{Z}=2.25\%$ the output curve for layer 1 drops below the curve for layer 0. Since $\braket{M^{l}_{\mathrm{QCNN}}}$ depends on the error rate on the lattice at layer $l$, this indicates that the density of errors has increased with the first pooling layer. Such behavior is expected as the error rate is close to the threshold. In this case, the number of errors removed by the first pooling layer is smaller than the reduction in the total number of qubits and the errors can only be cleared sufficiently with the subsequent pooling layer. This behavior also starts to become apparent in the second pooling layer for larger error rates and for our last test with $p_{X}=p_{Z}=10\%$ the error density after the second pooling layer even increases beyond the density after the first layer.
The phase recognition continues to agree with the ground states in the absence of Pauli-noise up to $p_{X}=p_{Z}=3\%$, where the input states up to the phase transition are more likely to be characterized correctly since the average QCNN output $\braket{\mathrm{QCNN}}$ rests above $50\%$ in the topological phase. In our tests, we find topological input states that are falsely characterized as trivial for $p_{j}>3\%(j=X,Z)$ and for strong Pauli noise at $p_{X}=p_{Z}=10\%$ no input state ensemble is characterized as topological. This behavior is expected as the correlation length of the syndromes of the Pauli errors becomes so long compared to the error correction capabilities of the QCNN, that the topological character of the input states remains masked.
[…]
@@
In conclusion, we generalized QCNNs to two-dimensional systems and showed that they recognize the topological phase of the Toric Code. The QCNN output is robust against errors below a threshold probability allowing for their realization on current quantum computers under NISQ conditions. Furthermore, all pooling layers can be efficiently performed in classical post-processing to facilitate the QCNN implementation with restricted circuit depths on NISQ devices. In future research, one can further investigate the quantum error correction procedure in the pooling layers to increase the error threshold and enhance the phase detection by a faster convergence to a step function with the number of pooling layers. The implementation of the convolutional layer as a quantum circuit opens the way for the efficient processing of quantum data with reduced sample complexity [35, 34]. We expect that our work motivates the exploitation of QCNNs for the characterization of less-understood quantum phases of matter such as symmetry-enriched topological order [43] and quantum spin liquids [18]. Implementing pooling layers as quantum circuits is an interesting possibility for efficiently characterizing such complex systems allowing for the parameterization and training of the QCNNs [32, 33, 44].
@@ A.1 Explicit Structure of the QCNN
In this section, we explain the explicit structure of our QCNN in full detail. For the convolution, the starting point is the generation of the reference state. The intuitive choice for the reference state is the ground state of the unperturbed Toric Code, $H_{\text{TC}}$. From [23] we know that there is an exact circuit that can prepare the intrinsic topological order of the Toric Code from an all-zero initialization of the qubits. This is achieved via the preparation of the plaquette stabilizers from a set of representative qubits. Specifically for one plaquette, we bring the representative qubits to the $\ket{+}$ state with a Hadamard gate and subsequently entangle the other qubits with CNOT gates that are controlled on the representative qubit. This is sufficient to implement the X-parity between the four qubits of each plaquette.
The representative qubits need to be chosen in a sequence that ensures that they have not been entangled with other qubits before. Throughout the bulk of the lattice, this method is easy to apply via the preparation of plaquettes in sequential columns. Only at the periodic boundaries it is important to be mindful of choosing the correct representative qubits, that have not yet been entangled by the first column of plaquettes. As neighboring vertex and plaquette stabilizers always share two qubits, this circuit also satisfies the requirement of the Z-parity between the qubits of each vertex stabilizer and therefore, the Toric Code ground state is correctly prepared. We will call this circuit $U_{\mathrm{prep}}$.
Accordingly, with $U^{\dagger}_{\mathrm{prep}}\ket{\psi_{\mathrm{toric}}}=\ket{0}^{\otimes N}$ we know that the inverse preparation circuit $U^{\dagger}_{\mathrm{prep}}$ takes the Toric Code ground state back to the all-zero state. If we apply the inverse preparation circuit $U^{\dagger}_{\mathrm{prep}}$ to a perturbed lattice state |$\ket{\phi}\neq\ket{\psi_{toric}}$, the circuit will map the parity of all plaquettes to the representative qubits.
In Figure 7 we show the full sequence of gates for the convolution $U_{\mathrm{C}}^{(0)}$ as in Equation 2 for a small $3\times 3$ example. This includes additional CNOT and SWAP gates that we need to add to $U^{\dagger}_{\mathrm{prep}}$ in order to also map the vertex stabilizers correctly to a shifted set of representative qubits. During this procedure, we are removing the mapping of the logical operators. Hence, we are losing all information about which state within the subspace of the degenerate ground states of the Toric Code was put in and the QCNN accordingly only preserves information about the topological phase.
@@
Figure 7: Sequence of all gates in the convolution $U_{\mathrm{C}}^{(0)}$ (see main text) for a $3\times 3$ lattice with periodic boundaries. Figures a) to e) depict the inverse Toric Code ground state preparation algorithm $U^{\dagger}_{\mathrm{prep}}$. The blue gates are CNOTs, for which the $+$ mark the respective targets and the circles mark the controlling qubits. Here we show how for the rightmost column of the lattice the representative qubits are the lower gray qubits, whereas for all other columns of plaquettes, the right green qubit is the representative qubit. This is an example of a set of representative qubits that ensures during the ground state preparation circuit $U_{\mathrm{C}}$ that the representative qubits have not yet been entangled before their plaquette is prepared. On lattice e) we also visualize the Hadamard gates (red), which also are part of the inverse preparation circuit. Figures f) to h) show additional CNOTs that are necessary to map the vertex stabilizers correctly. These gates do not act on the mapping of the plaquette stabilizers as they are only controlled on gray qubits on the horizontal lattice edges. In g) and h) we additionally implement SWAP gates over the periodic lattice edges to correct the broken translational symmetry in a) and b) from choosing gray qubits as representative qubits. Lastly, in figures h) to j) we reset the bottom left and top right qubits, before mapping the product of all other stabilizers of the corresponding basis to the qubits. Before the reset, the procedure of a) to g) mapped a logical X and a logical Z operator to these two qubits respectively. In figures i) and j) specifically, we apply CNOT gates, that target the qubit in the corner and are controlled by all other qubits of the corresponding sublattice. With the last step we remove all information about the logical subspace, but ensure that all stabilizer generators are mapped to the correct sites and thus can apply the local pooling operations in a translationally invariant way.
@@
After the implementation of the convolution, we now outline the structure of the pooling layer. The advantage of mapping the stabilizer measurements to individual qubits is that perturbations on the Toric Code can be easily revealed. Single-qubit perturbations induce error syndromes on the lattice of the Toric Code in its error-correcting functionality. As every qubit is part of four stabilizer elements, the single-qubit error will induce a negative parity on the neighboring stabilizer elements that anticommute with the error. In our case, such a single-qubit error will after the convolution give rise to a pair of Pauli-X errors on neighboring qubits. We show this property in Figure 8.
@@
Hence, the challenge is to implement a scheme of error correction that is suited to these characteristics. For the layered structure of the QCNN with the concept of renormalization, we want to choose a set of target qubits that are error-corrected. In general, we want to remove short-range behavior and evaluate only global topological order on the finally remaining output qubits. For error correction, one needs to find a procedure that is capable of removing any singular error and reducing the order of larger error clusters. A singular error here means that it is an error that is sufficiently far away from all other errors such that they do not interact, whereas an error cluster can mean directly neighboring or overlapping error patterns.
In this work, we propose to choose the set of target qubits $\{t_{j}\}$ to be evenly spaced out on the lattice with a periodicity of three. This ensures that each target qubit $t_{j}$ has a cross-shaped set of neighbors that are at the same time not nearest neighbors of another target qubit. Therefore, singular errors that are detected on these neighboring qubits can be unambiguously associated with the specific target qubit. After the convolution, this allows for the correction of all error pairs that originate from singular errors on the input and are located on the target qubit via the application of CNOT gates controlled on the four nearest neighbors of the target qubit. We call these the control qubits $\{c_{k}\}=\mathcal{N}(t_{j})$ with $\mathcal{N}(t_{j})$ referring to the neighborhood of $t_{j}$. However, the CNOTs will also propagate errors to the target qubits from error pairs that are on the qubits in between the target qubits but not on the targets themselves. Therefore, it is necessary to apply Toffoli gates in an additional step to the target qubits that are controlled by the control qubits $\{c_{k}\}$ and by their respective nearest neighbors without the target qubit $\{n_{l}\}=\mathcal{N}(c_{k})\backslash t_{j}$. This will then remove a wrongful correction that was introduced by the CNOT gates. According to the chosen set of target qubits $\{t_{j}\}$ we can calculate the required number of qubits in each layer $l$ for a specific depth $d$ of the QCNN as
$N_{\mathrm{qubits}}(l)=2\times 3^{2(d-l)}\,.$ (8)
The pooling is shown for a small example in Figure 9 and can be expressed as a unitary as
$U_{P}=\prod_{c\in\mathcal{N}(t)}\prod_{n\in\mathcal{N}(c)\backslash t}\textrm{Toff}_{nct}\textrm{CNOT}_{ct}\,,$ (9)
$U_{P}^{\dagger}Z_{t}U_{P}=$ $\left(\prod_{c\in\mathcal{N}(t)}\prod_{k\in\mathcal{N}(c)\backslash t}\textrm{Toff}_{kct}\textrm{CNOT}_{ct}\right)Z_{t}\left(\prod_{d\in\mathcal{N}(t)}\prod_{l\in\mathcal{N}(c)\backslash t}\textrm{CNOT}_{dt}\textrm{Toff}_{ldt}\right)=$
@@
$=$ $\left(\prod_{c\in\mathcal{N}(t)}Z_{c}\right)Z_{t}\left(\prod_{\begin{subarray}{c}c\in\mathcal{N}(t),\\ k\in\mathcal{N}(c)\backslash t\end{subarray}}\textrm{CZ}_{kc}\right)$
with the identities
$\textrm{CNOT}_{ct}Z_{t}\textrm{CNOT}_{ct}$ $=Z_{c}Z_{t}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{and}$
$\textrm{Toff}_{kct}Z_{t}\textrm{Toff}_{kct}$ =(|$=(\ket{11}\bra{11}_{kc}\otimes X_{t}+(1-\ket{11}\bra{11}_{kc})\otimes\mathds{1}_{t})Z_{t}\textrm{Toff}_{kct}$
@@
$=Z_{t}(1-2\ket{11}\bra{11})$
$=Z_{t}\textrm{CZ}_{kc}\,.$
In general, this pooling procedure as part of the QCNN can be written and implemented as a quantum circuit. However, as it only consists of controlled Pauli gates that act on Pauli strings followed by measurements, one can also perform these operations as classical post-processing on measurement snapshots after having performed the convolutional layer. Moreover, X-stabilizers and Z-stabilizers are separately processed in the pooling procedure. As a result, the convolutional layer can be replaced by the direct sampling of the ground states in both the Z-basis and X-basis. For each measurement snapshot in the Z-basis (X-basis), we can determine all Z-stabilizer (X-stabilizer) values. In comparison to the measurement after the convolution, the direct measurement of the input state in both Z-basis and X-basis increases sample complexity by a factor of two.
@@
Figure 9: Graphical depiction of gates in the pooling procedure. a) Green and gray qubits form the two shifted sublattices of the Toric Code. Encircled are the target qubits, that will be propagated to the higher layer. The red circles mark the target qubits for the sublattice of the qubits on vertical edges (pooling in the X-basis), whereas the blue circle marks the target qubits on horizontal lattice edges (pooling in the Z-basis). On the target qubits, we apply CNOT gates controlled by their nearest neighbors on their respective sublattices. For visibility reasons, we only show the CNOT gates for one target qubit of each sublattice. b) Here, we show only the sublattice of qubits on vertical lattice edges. Subsequently to the CNOT gates in a), we apply Toffoli gates controlled on the nearest and next-nearest neighbors of the target qubits. This ensures that pairs of errors that are anchored on the nearest neighbors but not on the target qubit do not cause a wrong flip of the target qubit. Specifically, such an error pair first triggers a flip of the target qubit via the CNOTs, which is then consequently removed after the application of the Toffolis.
A.2 Additional Details on the Numeric Simulations
In this section, we provide additional information on our implementation of the MPS. We use the TenPy python library [45] to initialize the MPS on an infinite cylinder geometry. Accordingly, the horizontal dimension $l_{1}$ is infinite with open boundaries and the vertical dimension $l_{2}$ is periodic. As a strategy to generate the MPS we run the infinite Density Matrix Renormalization Group (iDMRG) algorithm [41] in two distinct steps in order to consistently select a specific ground state of the degenerate manifold. In the first step, we add an energy penalty for the logical operators to select the +1 eigenstate of the Wilson and t’Hooft loop operators that span the periodic dimension of the cylindric lattice before we optimize without the penalty in a second iteration of iDMRG. This ensures that we avoid convergence to superpositions of the different ground states. We can then numerically compute ground states with a bond dimension of up to $\chi=2000$. However, we restrict the bond dimension to $\chi=1250$ for most plots in this manuscript, as we found that increasing it further did not lead to a visible change in our findings. Special attention is also required for the MPS initialization in the case of $h_{X}=h_{Z}$ as in Figure 4 b). As the ground state under this condition is two-fold degenerate [42], we first run iDMRG with $h_{X}=h-\delta$ and $h_{Z}=h+\delta$ for a small $\delta=0.05$ in order to select the corresponding eigenstate. In a second run of DMRG, we then optimize the MPS for $h_{X}=h_{Z}$ to converge to the correct ground state. Note that, for the other ground state obtained with $\delta=-0.05$, the QCNN output is similar to Figure 10 but the values in X- and Z-basis interchange.
Figure 10: QCNN output for the multicritical point along $h_{X}=h_{Z}$. In comparison to Figure 4 b) in the main text, we plot each pooling layer individually instead of taking the product over the pooling in both bases. We observe a splitting in the values of corresponding $d_{X}$ and $d_{Z}$. This results from the property that the ground state of the Toric Code manifold in this case is two-fold degenerate with a respective bias towards one type of stabilizer. If we initialize the MPS with $\delta=-0.05$, the behavior of both pooling bases will interchange.
As we discussed in the previous section, the QCNN can be entirely run as a quantum circuit but has the benefit that it can also be fully realized in classical post-processing. We sample spin configurations in the Z-basis and in the X-basis from the ground states. We determine the QCNN output in classical post-processing as explained in the previous section. For our simulations, we draw samples for a total of 486 qubits on a $2l_{1}l_{2}=2\times 27\times 9$ qubit lattice. A reduced schematic of the lattice can be found in Figure 11. In the lattice dimensions, the factor of two refers to the two shifted sublattices for qubits either on horizontal or vertical lattice edges. Furthermore, $l_{1}$ is the number of qubits in the horizontal dimension and $l_{2}$ in the vertical direction.
We chose these specific dimensions $l_{1}$ and $l_{2}$ such that we can apply two consecutive pooling layers, which fixes the vertical (periodic) dimension to $l_{2}=9$ qubits. Intuitively, the same minimum number of qubits should also suffice for the QCNN in horizontal dimension, but as the MPS is non-periodic in this direction, we have to use additional qubits as a buffer. This is necessary because the open boundaries break the properties of the convolution $U_{C}^{(0)}$ at the edges. We solve this problem by first applying the QCNN to all sampled qubits, but only extracting information from the qubits, for which the light cone is restricted to the bulk of the lattice and does not span out to the edges as sketched in Figure 11. This requirement must be respected in every layer of the QCNN. Hence, we can find a lower limit for the required number of qubits by defining the lattice configuration for the highest layer. From the QCNN structure, we know that the number of qubits increases by a factor of 9 for every pooling layer and the highest output layer consists of a column of two qubits. To mitigate the edge interactions on the non-periodic dimension of the cylinder code, we need to trim at least one column of qubits on each side of the contributing qubits in horizontal dimension. Accordingly, the output layer must be grown to a number of $N=3\times 2$ qubits to the triple of the original size in dimension $l_{1}$. As all qubits in the higher layers must be propagated from lower layers, the lower layers must all also be tripled in horizontal dimension to a final size of $l_{1}=27$, but all target qubits but the left- and rightmost column can contribute to the results.

### v2

@@ Abstract
with $K_{l}\in\{\sqrt{p_{\mathds{1}}}\mathds{1},\,\sqrt{p_{X}}X,\,\sqrt{p_{Y}}Y,\sqrt{p_{Z}}Z\}^{\otimes N}$ to the input qubits, where $p_{E}$ are probabilities of Pauli errors $E=X,Y,Z$ and $p_{\mathds{1}}=1-p_{X}-p_{Y}-p_{Z}$. As the QCNN only consists of Clifford and controlled Pauli gates, we can then subsequently track the evolution of the Pauli error strings on the lattice under the operations of the QCNN. It is sufficient to employ the following gate identities $X_{j}=H_{j}Z_{j}H_{j}$, $\textrm{CNOT}_{i,j}X_{i}=X_{i}X_{j}\textrm{CNOT}_{i,j}$ and $\textrm{CNOT}_{i,j}Z_{j}=Z_{i}Z_{j}\textrm{CNOT}_{i,j}$, where $H_{j}$ is the Hadamard gate on qubit $j$ and $\textrm{CNOT}_{i,j}$ applies a Pauli $X$ on qubit $j$ conditioned on qubit $i$.
We present results for the QCNN output, Eq. 2, for cases where the input state is the reference state of the topological phase, i.e. a ground state of $H_{\mathrm{TC}}$ as in Eq. 4, perturbed by Pauli-Z noise ($p_{X}=p_{Y}=0$) in Fig. 4, which shows a transition in the QCNN output over the Pauli-Z error probability $p_{Z}$. The QCNN output converges to the ideal noise-free value $\braket{M^{d}_{\mathrm{QCNN}}}=1$ with the increasing number of pooling layers $d$ below the threshold error probability $p_{\rm th}=2.28\%$ (red vertical line). On the other hand, the output decreases to zero above this threshold. In the supplemental material SM , we also show results for simultaneous Pauli-Z and Pauli-X noise, and demonstrate that the QCNN detects the phase transition at the critical strength $h_{Z}^{*}$ of the magnetic field in the presence of noise.
Figure 4: QCNN output for incoherent Pauli noise perturbing the Toric Code ground state as in Eq. 11 with $p_{X}=p_{Y}=0$. Via tracking the evolution of Pauli strings on the lattice under the convolution and the pooling layers we can classically simulate the QCNN output for large system sizes. For this plot, we calculate the results for about 9.6 Million qubits, which allows us to employ more pooling layers compared to the MPS samples. As for the case of the magnetic field, we find a transition in the QCNN output that becomes more pronounced with increasing depth of the QCNN. The red vertical line marks the error threshold of $p_{\mathrm{th}}=2.28\%$.
We now discuss the key difference between the QCNN, and approaches based on direct measurements and snapshot-based processing as in Cong et al. (2024). On most quantum processors, only measurements in local Pauli bases can be directly performed. In this case, plaquette operators $A_{p}$ and star operators $B_{v}$ cannot be simultaneously measured as they require readout in different Pauli bases. As a result, such direct measurements do not capture the correlations between these operators and only uncorrelated pooling, where the measurement outcomes of plaquette and star operators are processed separately, can be performed. In our QCNN, in contrast, all $A_{p}$ and $B_{v}$ operators are simultaneously measured after having applied the convolutional layer as a quantum circuit, see Fig. 2. Therefore, we can implement correlated pooling that takes into account correlations between these operators similarly to correlated quantum error correction Fowler (2013). See the supplemental material SM for details about correlated pooling and its numerical simulation.
Figure 5: Output of (a) uncorrelated pooling and (b) correlated pooling for the magnetic field strength $h_{Y}$ in the presence of Pauli-Y noise with $p_{Y}=3\%$ and for different depths $d$. Each data point represents 2000 samples from the MPS representation of the corresponding ground state with (a) the bond dimension $\chi=1000$ and the periodic dimension $l_{2}=9$, and (b) the bond dimension $\chi=1200$ and the periodic dimension $l_{2}=6$. This comparison shows that the correlated pooling scheme can precisely detect the first-order phase transition (vertical red line), whereas topological states close the phase transition are misclassified for uncorrelated pooling, see the inset.
To demonstrate the advantage of our QCNN, we consider incoherent Pauli-$Y$ errors that simultaneously flip the parity of two plaquette and two star operators, introducing correlations between them. The error threshold for Pauli-Y errors is considerably increased to $3.48\%$ for correlated pooling compared to $2.28\%$ for uncorrelated pooling, see the supplemental material SM . This shows that the error correction capability is boosted by taking correlations into account.
In addition to incoherent Pauli-Y errors, we assume a coherent magnetic field in the $Y$ direction that drives the Toric Code across a first-order phase transition into the trivial phase Dusuel et al. (2011). We plot the output for uncorrelated and correlated pooling in Figures 5a and 5b for an error probability $p_{Y}=3\%$. The output of the QCNN with correlated pooling is larger for all states in the topological phase than the output for uncorrelated pooling. Due to the limited ability to correct correlated errors, uncorrelated pooling misclassifies topological states close to the phase boundary as the output for $d=1,2$ is smaller than for $d=0$, see the inset of Fig. 5a. In contrast, the QCNN with correlated pooling correctly recognizes all topological states. Hence, by exploiting correlations between measurement outcomes, the QCNN correctly recognizes the topological phase in the presence of strong correlated errors close to the phase transition, where classical approaches misclassify topological states.
Finally, we test the QCNN on ground states of $H_{\textrm{mag}}$ with magnetic field simultaneously acting in both X- and Z-direction. First, we consider constant $h_{X}=0.1$ leading to a non-vanishing probability of error syndromes $B_{i}=-1$ on vertices and thus $M_{\mathrm{QCNN}}^{0}<1$, see Fig. 6a. Second, we consider the transition into the trivial phase driven by perturbations in both X- and Z-directions for $h_{X}=h_{Z}=h$ Vidal et al. (2009). The QCNN output rapidly decreases with increasing depth $d$ above the multicritical point $h=0.35$, see Fig. 6b. This provides clear evidence that the QCNN recognizes the phase transition also for these cases.
Figure 6: a) Output of the QCNN for the MPS samples under a variation of $h_{Z}$ in the presence of a weak magnetic field in X-direction with $h_{X}=0.1$ with the bond dimension $\chi=1250$. The output is displayed for simultaneous pooling in the X(Z)-basis for the corresponding QCNN depth $d_{X}$($d_{Z}$). The phase transition for $h_{Z}$ is correctly recognized in the presence of a field in X-direction. Furthermore, the output of the pooling layers $d_{Z}$ quickly converges to unity due to the weak $h_{X}=0.1$. b) Detection of the multicritical point Vidal et al. (2009) in the presence of X- and Z-field with $h=h_{X}=h_{Z}$ for the bond dimension $\chi=1500$. We take the product of the pooling in the X- and Z-basis for each QCNN layer to show the combined phase recognition along the multicritical line $h_{X}=h_{Z}$. A split plot can be found in the supplemental material SM .
@@
In conclusion, we generalized QCNNs to two-dimensional systems and showed that they recognize the topological phase of the Toric Code. The QCNN output is robust against errors below a threshold probability allowing for their realization on current quantum computers under NISQ conditions. In processing quantum states via the QCNN quantum circuit, correlations between all stabilizer elements of the Toric Code are taken into account. As a result, the threshold probability for correlated errors is increased compared to classical approaches based on direct measurements as in Cong et al. (2024). Therefore, the QCNN correctly identifies the topological phase in the presence of strong correlated errors, where classical approaches misclassify topological states. Furthermore, all pooling layers can be efficiently performed in classical post-processing to facilitate the QCNN implementation with restricted circuit depths. In future research, one can further investigate the error correction procedure in the pooling layers to increase the error threshold and enhance the phase detection by a faster convergence to a step function with the number of pooling layers. We expect that our work motivates the exploitation of QCNNs for the characterization of less-understood quantum phases of matter such as symmetry-enriched topological order Mesaros and Ran (2013) and quantum spin liquids Savary and Balents (2016). Implementing the pooling layers as quantum circuits is an interesting possibility for characterizing such complex systems with reduced sample complexity, and allowing for the parametrization and training of the QCNNs Caro et al. (2022); Liu et al. (2022); Pesah et al. (2021).
@@ A.1 Explicit Structure of the QCNN
In this section, we explain the explicit structure of our QCNN in full detail. We start by discussing the overall structure of the QCNN. The QCNN is a phase recognition circuit, which is an analog of classical Convolutional Neural Networks (CNNs). The QCNN can be written as a unitary of the form Cong et al. (2019); Herrmann et al. (2022); Lake et al. (2022); Zapletal et al. (2023a),
$U_{\mathrm{QCNN}}=U_{\mathrm{FC}}U_{\mathrm{CP}}^{(d)}...U_{\mathrm{CP}}^{(2)}U_{\mathrm{CP}}^{(1)}\,$, (7)
where the unitary $U_{\mathrm{CP}}^{(j)}$ represents a combination of a convolutional and a pooling layer, $d$ is the number of subsequent layers and the unitary $U_{\mathrm{FC}}$ is a fully connected layer that prepares the state of the output qubits, which are subsequently measured. For the $j^{\mathrm{th}}$ layer in the QCNN structure, we specifically define the unitary $U_{\mathrm{CP}}^{(j)}$ as $U_{\mathrm{CP}}^{(j)}=\tilde{U}_{\mathrm{P}}^{(j)}(U_{\mathrm{C}}^{(j)})^{\dagger}U_{\mathrm{C}}^{(j-1)}$, where $\tilde{U}_{\mathrm{P}}^{(j)}$ is the unitary implemented by the pooling procedure, $(U_{\mathrm{C}}^{(j)})^{\dagger}$ is the adjoint of the convolution of the current layer and $U_{\mathrm{C}}^{(j-1)}$ is the convolution of the preceding layer j1j-1.
If we construct the convolutional layers from Clifford gates and the pooling procedure via single qubit Pauli-operations that are controlled by the results of measurements of other qubits, we can shorten the QCNN circuit dramatically by pushing $U_{\mathrm{C}}^{(j)}$ through the pooling circuit $\tilde{U}_{\mathrm{P}}^{(j)}$ Herrmann et al. (2022); Zapletal et al. (2023b). We then obtain a modified pooling layer $U_{\mathrm{P}}^{(j)}$ and subsequently cancel all convolutions except the first one, such that we can write the QCNN unitary as
$U_{\mathrm{QCNN}}=U_{\mathrm{FC}}U_{\mathrm{P}}^{(d)}...U_{\mathrm{P}}^{(2)}U_{\mathrm{P}}^{(1)}U_{\mathrm{C}}^{(0)}\,.$ (8)
As a consequence, only the initial convolutional layer $U_{\mathrm{C}}^{(0)}$ is executed on the quantum device, while all other layers can be computed in classical post-processing. We define $U_{C}=U_{\mathrm{C}}^{(0)}$ for brevity.
@@
The representative qubits need to be chosen in a sequence that ensures that they have not been entangled with other qubits before. Throughout the bulk of the lattice, this method is easy to apply via the preparation of plaquettes in sequential columns. Only at the periodic boundaries it is important to be mindful of choosing the correct representative qubits, that have not yet been entangled by the first column of plaquettes. As neighboring vertex and plaquette stabilizers always share two qubits, this circuit also satisfies the requirement of the Z-parity between the qubits of each vertex stabilizer and therefore, the Toric Code ground state is correctly prepared. We will call this circuit $U_{\mathrm{prep}}$, see Fig. 7.
@@
In Fig. 7 we show the full sequence of gates for the convolution $U_{\mathrm{C}}^{(0)}$ as in the main text for a small $3\times 3$ example. This includes additional CNOT and SWAP gates that we need to add to $U^{\dagger}_{\mathrm{prep}}$ in order to also map vertex stabilizers to representative qubits, as we have already done with the plaquettes. During this procedure, we are also removing the mapping of the logical operators by resetting two qubits, see Fig. 7h. Hence, we are losing all information about which state within the subspace of the degenerate ground states of the Toric Code was put in and the QCNN accordingly only preserves information about the topological phase.
@@
Figure 7: Sequence of all gates in the convolution $U_{\mathrm{C}}^{(0)}$ (see main text) for a $3\times 3$ lattice with periodic boundaries. Figures a) to e) depict the inverse Toric Code ground state preparation algorithm $U^{\dagger}_{\mathrm{prep}}$. The blue gates are CNOTs, for which the ++ mark the respective targets and the circles mark the controlling qubits. Here we show how for the rightmost column of the lattice the representative qubits are the lower gray qubits, whereas for all other columns of plaquettes, the right green qubit is the representative qubit. This is an example of a set of representative qubits that ensures during the ground state preparation circuit $U_{\mathrm{C}}$ that the representative qubits have not yet been entangled before their plaquette is prepared. On lattice e) we also visualize the Hadamard gates (red), which also are part of the inverse preparation circuit. Figures f) to h) show additional CNOTs that are necessary to map the vertex stabilizers correctly. These gates do not act on the mapping of the plaquette stabilizers as they are only controlled on gray qubits on the horizontal lattice edges. In g) and h) we additionally implement SWAP gates, denoted by the cross symbol $\times$, over the periodic lattice edges to correct the broken translational symmetry in a) and b) from choosing gray qubits as representative qubits. Lastly, in figures h) to j) we reset the bottom left and top right qubits, before mapping the product of all other stabilizers of the corresponding basis to the qubits. Before the reset, the procedure of a) to g) mapped a logical X and a logical Z operator to these two qubits, respectively. In figures i) and j) specifically, we apply CNOT gates that target the qubit in the corner and are controlled by all other qubits of the corresponding sublattice. With this step we remove all information about the logical subspace, but ensure that all stabilizer generators are mapped to the correct sites and thus can apply the local pooling operations in a translationally invariant way. Lastly, in figures k) and l), we swap each column of qubits on horizontal lattice edges one step to the right, such that syndromes of Pauli-Y errors on the lattice create errors on four qubits in a diamond shape. This is only necessary as a preparation for the pooling scheme if there are correlated errors that affect both horizontal and vertical lattice edges simultaneously.
@@
After the implementation of the convolution, we now outline the structure of the pooling layer. The advantage of mapping the stabilizer measurements to individual qubits is that perturbations on the Toric Code can be easily revealed. Single-qubit perturbations induce error syndromes on the lattice of the Toric Code in its error-correcting functionality. As every qubit is part of four stabilizer elements, a single-qubit Pauli error will induce a negative parity on the neighboring stabilizer elements that anticommute with the error. In our case, such a single-qubit error will be converted to a pair of Pauli-X errors on neighboring qubits after the convolution. We show this property in Fig. 8.
Hence, the challenge is to implement an error correction scheme that exploits the correlated error pairs generated by the convolution $U_{\mathrm{C}}^{(0)}$. We choose a set of target qubits $\{t_{j}\}$ that are error-corrected and propagated to the next layer. In general, we want to remove short-range behavior and evaluate only global topological order on the finally remaining output qubits. As in quantum error correction, one needs to find a procedure that is capable of removing any singular error and reducing the size of error clusters relative to the total number of qubits. By singular error, we here mean that it is an error that is sufficiently far away from all other errors such that they do not interact. On the other hand, an error cluster can mean directly neighboring or overlapping error patterns after the convolutional layer.
@@
$N_{\mathrm{qubits}}(l)=2\times 3^{2(d-l)}\,.$ (9)
The pooling is shown for a small example in Fig. 10 and can be expressed as a unitary as
$U_{P}=\prod_{c\in\mathcal{N}(t)}\left(\prod_{n\in\mathcal{N}(c)\backslash t}\textrm{Toff}_{nct}\right)\textrm{CNOT}_{ct}\,$, (10)
@@
where we have used the identities
$\textrm{CNOT}_{ct}Z_{t}\textrm{CNOT}_{ct}$ $=Z_{c}Z_{t}\penalty 10000\ \penalty 10000\ \mathrm{and}$